How do you solve 3 over the square root of 48
Answer by checkley71(8403)

3/SQRT48 OR 3/SQRT(4*4*3) [NOTE 4*4*3=48] THEREFORE THE SQRT OF 4*4*3 IS SQRT16*3 THE SQRT OF 16=4 THUS THE SQRT16*3=4SQRT3
THUS THE ANSWER IE 3/4SQRT3